# Frequentist Validity of Epistemic Uncertainty Estimators

### v1

@@ 5.1 Uncertainty Decomposition
MI as a measure of information gain has demonstrated success in active learning (Houlsby et al.,, 2011; Gal et al.,, 2017), and the decomposition we consider has demonstrated success over others in OOD and misclassification detection tasks (Kotelevskii et al.,, 2025). Beyond the specific information theoretic uncertainty decomposition considered in our work, a variety of other decompositions have also been proposed (Senge et al.,, 2014; Kendall and Gal,, 2017; Hofman et al.,, 2024; Lahlou et al.,, 2023; Schweighofer et al.,, 2023; Smith et al.,, 2025) and MI does have drawbacks (Wimmer et al.,, 2023). Indeed, epistemic and aleatoric uncertainty are often not cleanly separable and take on different meanings in different contexts (Kirchhof et al.,, 2025). Our work is not intended to propose the optimality of MI as an epistemic uncertainty assessment, but rather to simply study its properties from a frequentist perspective motivated by its widespread use, strong conceptual foundation and practical success.
@@ 5.2 Bayesian and Frequentist Equivalence
Rubin, (1981) considers bootstrap from a Bayesian lens, although this interpretation is hard to reconcile with traditional Bayesian inference since it considers priors over the data rather than the parameters. Closest to our paper is the work by Newton and Raftery, (1994) which studied the use of bootstrap as a way of generating samples from a Bayesian posterior. Our work expands upon this idea by specifically leveraging this connection to connect frequentist and Bayesian notions of uncertainty. Moreover, we operationalize this idea in a modern context with DNNs and use it to study deep ensembles. Finally, Fong et al., (2023) also connects Bayesian parameter uncertainty to frequentist uncertainty arising due to lack of data. While their work conceptually shares connections to ours, it consider different practical applications.
@@ 5.3 Deep Ensembles
Deep ensembles (Lakshminarayanan et al.,, 2017) have demonstrated strong uncertainty quantification performance over other popular uncertainty quantification methods (Ovadia et al.,, 2019; Ashukha et al.,, 2020). Several reasons behind their success have been put forth (Fort et al.,, 2020), including Bayesian ones (Wilson and Izmailov,, 2020) but rigorous frequentist perspectives are lacking. The use of bootstrap in the construction of deep ensembles has also been investigated but rejected because bagged ensembles were found to have worse predictive accuracy than ensembles over random seeds that use all data points (Lee et al.,, 2015). The use of Dirichlet weights instead of multinomial has never been investigated to the authors’ knowledge, which we believe to be empirically more suitable (Sec. 2.2).
@@ 6 Discussion
Limitations: Firstly, our estimator requires retraining the model which can be prohibitive in the modern era due to computational limitations and access restrictions. On the theoretical front, Theorem 2 only proves the asymptotic validity of our estimator. Since epistemic uncertainty fundamentally targets uncertainty due to lack of data, asymptotic guarantees can appear meaningless. To alleviate this concern, we show the success of our in a low data regime—active learning. Our theorems also rely on assumptions that can be unrealistic in practical settings. For example, correct specification of the model and identifiability of parameters would never be expected to hold for DNNs. While promising, our empirical investigations are limited and more comprehensive studies are needed to draw rigorous conclusions on practical validity and success.
@@
Future Directions: Our proof-of-concept experiment in Sec. 4.4 proposes the idea of connecting data attribution methods to uncertainty quantification. Indeed, the variance of the influence function is also the variance of the estimator (Tsiatis,, 2007). This open the doors to studying “uncertainty attribution”—identifying data points or modalities responsible for model uncertainty—by estimating their influence function. Our investigations in Sec. 4.3 that attribute uncertainty to training data variability and optimization stochasticity are an example of this idea and we hope to investigate this more deeply in a future work.
Acknowledgements
The authors thank Dr. Dan Kluger for his kind help and insights into the proofs. A.J. and S.B. were partly supported by the ARPA-H ADAPT program.
Appendix
@@ A.3 Theorem 1
Under the assumptions (A1)-(A7), we have the asymptotic expansion
@@ A.3.1 Proof
Under assumptions  (A1)(A7), we can apply the Bernstein-von Mises theorem (van der Vaart, (2000) Theorem 10.1). Note that Bernstein von-Mises implies convergence in total variation distance which directly implies convergence in distribution. Hence, we have
$\sqrt{n}\left(\theta-\theta_{0}\right)\big|\mathcal{D}_{n}\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,\mathcal{I}^{-1}(\theta_{0}))\qquad\textrm{in $P_{\theta_{0}}$-probability}.$
@@
$\sup_{h\in BL_{1}}\left|\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}\left[h\left(\sqrt{n}(\theta-\theta_{0}\right)\right]-\mathbb{E}_{Z\sim\mathcal{N}(0,\mathcal{I}^{-1}(\theta_{0}))}[h(Z)]\right|\stackrel{{\scriptstyle P_{\theta_{0}}}}{{\rightarrow}}0,$
where $BL_{1}$ is the space of functions $h:\mathbb{R}^{p}\mapsto\mathbb{R}$ with Lipschitz norm bounded by 1. Recall that if both expectations on the left hand side were unconditional, limit of the left hand side going to 0 is the standard definition of convergence in distribution. See van der Vaart, (2000) Sec. 23.2.1 for further discussion.
We shall henceforth continue with the $(.)\stackrel{{\scriptstyle d}}{{\rightarrow}}$ (.)  in $P_{\theta_{0}}$ probability” notation for ease of exposition.
Note that (A3) allows us to apply the functional delta method (see van der Vaart, (2000) chapter 20) to evaluate the limit of the function $\theta\mapsto\hat{p}_{k}(X_{\mathrm{test}};\theta)$ under the posterior. This yields,
$\sqrt{n}\left(\hat{p}_{k}(X_{\mathrm{test}};\theta)-\hat{p}_{k}(X_{\mathrm{test}};\theta_{0})\right)\big|\mathcal{D}_{n}\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,\sigma^{2}_{k})\qquad\textrm{in $P_{\theta_{0}}$-probability},$ (4)
where $\sigma^{2}_{k}=\left[\frac{\partial\hat{p}(X_{\mathrm{test}};\theta_{0})}{\partial\theta}^{T}\mathcal{I}^{-1}(\theta_{0})\frac{\partial\hat{p}(X_{\mathrm{test}};\theta_{0})}{\partial\theta}\right]_{k,k}$. Finally, (A7) implies uniform integrability allowing us to conclude that the limit of the moment is the moment of the limiting distribution:
$\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}[\hat{p}_{k}(X_{\mathrm{test}};\theta)]\stackrel{{\scriptstyle P_{\theta_{0}}}}{{\rightarrow}}\hat{p}_{k}(X_{\mathrm{test}};\theta_{0}),$
$\mathbb{V}\textrm{ar}_{\theta\sim p(\theta|\mathcal{D}_{n})}\left[\sqrt{n}\hat{p}_{k}(X_{\mathrm{test}};\theta)\right]\stackrel{{\scriptstyle P_{\theta_{0}}}}{{\rightarrow}}\sigma_{k}^{2}.$
@@
$\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}\left[n\frac{(\hat{p}_{k}(X_{\mathrm{test}};\theta)-\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}[\hat{p}_{k}(X_{\mathrm{test}};\theta)])^{3}}{\hat{p_{k}^{\prime 2}}}\right]\stackrel{{\scriptstyle P_{\theta_{0}}}}{{\rightarrow}}0.$
@@
$n\mathrm{I}(Y_{\mathrm{test}};\theta|X_{\mathrm{test}},\mathcal{D}_{n})\stackrel{{\scriptstyle P_{\theta_{0}}}}{{\rightarrow}}\frac{1}{2}\sum_{k=1}^{K}\frac{\sigma_{k}^{2}}{\hat{p}_{k}(X_{\mathrm{test}};\theta_{0})}.$
@@ A.4 Theorem 2
Under the assumptions (A1)-(A12),
$\frac{\mathrm{I}(Y_{\mathrm{test}};\theta|X_{\mathrm{test}},\mathcal{D}_{n})}{\mathrm{I}_{b}(X_{\mathrm{test}},\mathcal{D}_{n})}\stackrel{{\scriptstyle P_{\theta_{0}}}}{{\rightarrow}}1\qquad\forall\ X_{\mathrm{test}},$
@@ A.4.1 Proof
where $\hat{p}_{k}^{\prime}$ is now some value between $\hat{p}_{k}(X_{\mathrm{test}};\hat{\theta}_{b})$ and $\mathbb{E}_{\hat{\theta}_{b}\sim p_{b}(\hat{\theta}_{b}|\mathcal{D}_{n})}[\hat{p}_{k}(X_{\mathrm{test}};\hat{\theta}_{b})]$ for k=1,,Kk=1,...,K. Next, from Theorem 10.16 of Kosorok, (2007), we have
$\sqrt{n}\left(\hat{\theta}_{b}-\hat{\theta}\big|\mathcal{D}_{n}\right)\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,\mathcal{I}^{-1}(\theta_{0}))\qquad\textrm{in $P_{\theta_{0}}$-probability}$
@@
We can now apply the delta method for bootstrap (van der Vaart, (2000) Theorem 23.9) to this and proceed exactly as in the Bayesian case to conclude
$n\mathrm{I}_{b}(X_{\mathrm{test}},\mathcal{D}_{n})\stackrel{{\scriptstyle P_{\theta_{0}}}}{{\rightarrow}}\frac{1}{2}\sum_{k=1}^{K}\frac{\sigma_{k}^{2}}{\hat{p}_{k}(X_{\mathrm{test}};\theta_{0})}.$
@@
$\frac{n\mathrm{I}(Y_{\mathrm{test}};\theta|X_{\mathrm{test}},\mathcal{D}_{n})}{n\mathrm{I}_{b}(X_{\mathrm{test}},\mathcal{D}_{n})}\stackrel{{\scriptstyle P_{\theta_{0}}}}{{\rightarrow}}1.$
@@ Appendix B Additional Experiments and Experimental Details
B.1 Validation of the Plug-In Principle
Experimental Details: The original SVHN (Netzer et al.,, 2011) dataset contains 73257 training set images, 26032 test set images and 531131 extra images. We combine the original training fold and extra images fold and then use this joint set to create 11 training datasets of 50,000 images each for our experiment. We keep the test set the same.
@@ B.1 Validation of the Plug-In Principle
We use a standard ResNet-18 (He et al.,, 2015) architecture but modify the first convolutional layer to have a kernel size of 3, stride of 1, padding of 1 to allow the network to work with 32 $\times$32 size images. Further, we modify the last fully connected later to only have 10 output neurons for a 10 class classification task. We use cross entropy loss and Adam (Kingma and Ba,, 2015) as an optimizer with learning rate of 0.001 and betas of 0.9 and 0.999. We use a batch size of 128 and train for 15 epochs.
The rest of the experiment proceeds as described in the main text.
B.2 Active Learning
Experimental Details: We follow the exact same setup as in Smith et al., (2023). We use the MNIST (LeCun et al.,, 1998) dataset. We first randomly select 4000 images per class to form the pool dataset. Out of these, we initialize the labeled dataset with 2 images per class and a validation dataset of 6 images per class. We keep the test set the same as in the original MNIST dataset.
We use a 2 layer convolutional network with 2 fully connected layers at the end. We use Dropout layers with probability $0.5$ between each convolutional and fully connected layer for all acquisition functions, but only for the Dropout estimate of mutual information, we keep the Dropout layers active during inference as in Gal and Ghahramani, (2016).
We train for up to 50,000 steps of full batch gradient descent with a learning rate of 0.01 and weight decay of 0.0001. We track validation log loss and after 5000 steps of non improvement, we trigger an early stopping routine which restores the model to the best validation set performance and terminates training.
The rest of the experiment proceeds as described in the main text.
B.3 Epistemic Uncertainty Decomposition
Experimental Details: We use the CIFAR-10 dataset (Krizhevsky,, 2009) with the original training and test set folds.
We use a standard ResNet-18 (He et al.,, 2015) architecture but modify the first convolutional layer to have a kernel size of 3, stride of 1, padding of 1 to allow the network to work with 32 $\times$32 size images. Further, we modify the last fully connected later to only have 10 output neurons for a 10 class classification task. We use cross entropy loss and Adam (Kingma and Ba,, 2015) as an optimizer with learning rate of 0.001 and betas of 0.9 and 0.999. We use a batch size of 128 and train for 15 epochs.
The rest of the experiment proceeds as described in the main text.
B.3.1 OOD Detection Results
We also assess the utility of our epistemic uncertainty estimates on a downstream out-of-distribution (OOD) detection task. We create 19 OOD versions of the CIFAR-10 test by applying the widely accepted corruptions proposed by Hendrycks and Dietterich, (2019) at 5 severity levels each. We then perform OOD detection on a test set created by appending the corrupted images to the original uncorrupted test set.
Table 1 lists the AUC-ROC values averaged across the 19 corruptions for each severity level for different uncertainty estimates. Specifically, $\mathrm{I}^{\textrm{deep ensemble}},\mathrm{I}^{\textrm{seeds}}_{b},\mathrm{I}^{\textrm{resampling}}_{b},\mathrm{I}_{b}$ are various mutual information estimates as described in the main text. $\mathrm{H}^{\textrm{deep ensemble}}$ is the entropy of the ensemble prediction. $\mathrm{H}_{b}$ is the entropy of the bootstrap ensemble prediction (i.e. average prediction across all bootstrap models). “Top Softmax Score” is a baseline that simply consists of using the highest class probability to order the points to form the ROC curve.
Firstly, note that $\mathrm{I}^{\textrm{resampling}}_{b}$ performs the worst perhaps surprisingly, and $\mathrm{I}^{\textrm{seeds}}_{b}$ and $\mathrm{I}_{b}$ perform very similarly as may be expected due to $\mathrm{I}^{\textrm{resampling}}_{b}$ being small and hence $\mathrm{I}_{b}\approx\mathrm{I}^{\textrm{seeds}}_{b}$. $\mathrm{I}^{\textrm{deep ensemble}}$ performs marginally better but all epistemic uncertainty measures perform worse than the simple baseline. The total uncertainty measures (based on predictive entropy) perform better with the predictive entropy of the ensemble being the best as may be expected in light of previous similar empirical findings. Furthermore, Kotelevskii et al., (2025) also observe that total entropy measures perform better for OOD detection tasks and offer discussions as to why.
Table 1: AUC-ROC for OOD detection of CIFAR-10 Corruptions
$\mathrm{I}^{\textrm{deep ensemble}}$ $\mathrm{I}^{\textrm{seeds}}_{b}$ $\mathrm{I}^{\textrm{resampling}}_{b}$ $\mathrm{I}_{b}$ $\mathrm{H}^{\textrm{deep ensemble}}$ $\mathrm{H}_{b}$ Top Softmax Score
Severity 1 0.565 0.561 0.557 0.561 0.566 0.562 0.565
Severity 2 0.613 0.606 0.599 0.606 0.615 0.608 0.613
Severity 3 0.649 0.640 0.630 0.640 0.652 0.643 0.649
Severity 4 0.682 0.674 0.661 0.673 0.687 0.678 0.683
Severity 5 0.724 0.716 0.703 0.716 0.735 0.725 0.728
B.4 Data Attribution

### v2

@@ 5.1 Uncertainty Decomposition
MI as a measure of epistemic uncertainty has demonstrated success in active learning (Houlsby et al., 2011; Gal et al., 2017), out-of-distribution and misclassification detection tasks (Kotelevskii et al., 2025). Beyond the specific information theoretic uncertainty decomposition considered in our work, a variety of other ways of quantifying epistemic uncertainty have also been proposed (Senge et al., 2014; Kendall and Gal, 2017; Hofman et al., 2024; Lahlou et al., 2023; Schweighofer et al., 2023; Smith et al., 2025) and MI does have drawbacks (Wimmer et al., 2023). See (Hüllermeier and Waegeman, 2021) for a comprehensive review of the topic. Indeed, epistemic and aleatoric uncertainty are often not cleanly separable and take on different meanings in different contexts (Kirchhof et al., April 28, 2025). We use MI for our investigations because of its widespread use, strong conceptual foundation and practical success; our work is not intended to address the optimality of MI as an epistemic uncertainty assessment.
Similar to our work, (Huang et al., 2023) and (Neal et al., 2019) also decompose the variance in model predictions for regression tasks into data variability and training randomness components. However, quite different to our contribution, (Huang et al., 2023) used this decomposition to study ways of theoretically eliminating training randomness using the NTK approximation (Jacot et al., 2018), while (Neal et al., 2019) used this decomposition to investigate the bias-variance tradeoff.
@@ 5.2 Bayesian and Frequentist Equivalence
(Rubin, 1981) considers bootstrap from a Bayesian lens, although this interpretation is hard to reconcile with traditional Bayesian inference since it considers priors over the data rather than the parameters. Closest to our theoretical results is the work by (Newton and Raftery, 1994) which studied the use of bootstrap as a way of generating samples from a Bayesian posterior. Our work expands upon this idea by specifically leveraging this connection to connect frequentist and Bayesian notions of uncertainty. Moreover, we operationalize this idea in a modern context with DNNs and use it to study deep ensembles. Finally, (Fong et al., 2023) also connects Bayesian parameter uncertainty to frequentist uncertainty arising due to lack of data. While their work conceptually shares connections to ours, it consider different practical applications.
@@ 5.3 Deep Ensembles
The uncertainty quantification capabilities of deep ensembles (Lakshminarayanan et al., 2017) were studied in depth by (Ovadia et al., 2019) and (Ashukha et al., 2020). Broadly speaking, these studies observed that deep ensembles have well calibrated probabilistic forecasts both in and out-of-domain, and exhibit strong performance on downstream tasks such as out-of-distribution detection and misclassification detection.
Several reasons behind their success have been put forth (Fort et al., 2020), including Bayesian ones (Wilson and Izmailov, 2020). These works primarily note that members of a deep ensemble capture different modes of the loss landscape and can be seen as high quality samples from the posterior. Our work is complementary to this perspective and provides a frequentist understanding of what uncertainties deep ensembles capture.
5.4 Bootstrapping for Uncertainties in Deep Learning
The idea of using bootstrap to estimate uncertainties in machine learning has been explored by several works in different contexts. (Jürgens et al., 2024) investigates the suitability of evidential deep learning (Sensoy et al., 2018) for uncertainty quantification by comparing them to the bootstrapped estimates of the model parameters, which they believe accurately capture uncertainty in model predictions due to randomness in the training data. (Jiménez et al., 2025) investigates variance based measures of uncertainty in a regression setting to also note that deep ensembles only capture uncertainty due to training variability which often dominates over uncertainty due to randomness in training data, similar to our findings in Sec. 3. Lastly, (Shaker and Hüllermeier, 2020) have also considered approximating the posterior distribution by the trees in a random forest, which can be seen as bootstrap estimates, for the purposes of mutual information estimation as in our work.
@@ 6 Discussion
6.1 Limitations
Theorem 3.2 only proves the asymptotic validity of the bootstrap estimator which can weaken the theoretical link between deep ensembles and Bayesian inference in low data regimes. Furthermore, our theorems also rely on assumptions that can be unrealistic in practical settings. In particular, correct specification of the model and identifiability of parameters can certainly appear tenuous for DNNs. Finally, while we aimed to cover a wide range of deep learning settings in our empirical studies, we recognize that more evaluation can always be carried out to lend robustness to our conclusions.
6.2 Future Directions
Figure 4: Influence Function Approximation: Estimates of epistemic uncertainty using Dropout, versus using influence functions to approximate predictions from bootstrapped datasets.
We identify an interesting link between the bootstrap estimator in Sec. 3.3 and data attribution methods such as influence functions (Hampel et al., 2011; Koh and Liang, 2017; Giordano et al., 2019) and datamodels (Ilyas et al., 2022; Park et al., 2023) which aim to identify training data points most responsible for a model’s predictions. These methods allow for estimation of predictions under perturbation of training data and can therefore be used to approximate the predictions of a model trained on bootstrapped data needed for our proposed measure. Any data attribution algorithm that yields a mapping from training data to model predictions can be leveraged to estimate the variance terms in (2), and hence give epistemic uncertainty estimates. We investigate the performance of such an MI estimate based on influence functions; see Appendix B.2 for a similar experiment with datamodels and hyperparameter details.
We train a single convolutional network called the LeNet-5 on MNIST data (LeCun et al., 1998). Next, we calculate the influence functions for each data point as described in (Koh and Liang, 2017) for the model parameters in the last two fully connected layers. For each data point, we multiply the influence function by the change in the weighting under bootstrap and add all these products to the parameters to obtain an approximation for model parameters under a bootstrap reweighting. We obtain predictions from 100 such approximated models and use Algorithm 1 to obtain MI estimates. To obtain a Bayesian estimate of MI, we follow the Bayesian interpretation of Monte Carlo Dropout proposed in (Gal and Ghahramani, 2016) wherein Dropout is enabled at inference time and stochastic forward passes are performed to obtain samples from the Bayesian predictive distribution.
Fig. 4 shows the influence function approximation and Dropout estimates. While there is correlation, influence function estimates are an order of magnitude smaller. We hypothesize that this is because the influence function approximation only uses a single trained model and hence fails to capture the randomness due to training stochasticity.
@@ 6.2 Future Directions
Nonetheless, our experiment serves as proof-of-concept for studying “uncertainty attribution”—identifying data points or modalities responsible for model uncertainty—by estimating their influence function. Since the variance of an estimator is equal to variance of its influence function (Tsiatis, 2007), we can note that training data points with high influence on a test prediction drive up the variance of the influence function and hence lead to higher epistemic uncertainty. We hope to explore this more in a future work.
7 Conclusion
Our paper aims to investigate the uncertainty quantification capabilities of deep ensembles from a frequentist perspective. Specifically, we study the epistemic uncertainty captured by deep ensembles, as measured by the mutual information between the model parameters and the data label. We derive an asymptotic expansion for this mutual information and use it to motivate the construction of an asymptotically correct bootstrap based estimator. Bayesian formulations cannot account for randomness in model parameter selection, while deep ensembles exclusively make use of this randomness. To bridge this gap, we make use of the bootstrap estimator—we show how the bootstrap estimate can be seen as an additive composition of the uncertainty from training randomness and uncertainty due to data sampling. We then conduct empirical studies which establish that training randomness dominates over data sampling, and that deep ensemble estimates of epistemic uncertainty closely capture the training randomness. Therefore, our investigations indicate that deep ensembles succeed at uncertainty quantification tasks by targeting the dominant contribution to an asymptotically correct estimate of the true Bayesian epistemic uncertainty.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.3 Theorem 1
Under the assumptions 1-7, we have the asymptotic expansion
@@ A.3.1 Proof
Under assumptions  17, we can apply the Bernstein-von Mises theorem ((van der Vaart, 2000) Theorem 10.1). Note that Bernstein von-Mises implies convergence in total variation distance which directly implies convergence in distribution. Hence, we have
$\sqrt{n}\left(\theta-\theta_{0}\right)\big|\mathcal{D}_{n}\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,\mathcal{I}^{-1}(\theta_{0}))\qquad\textrm{in $P_{\theta_{0}}^{n}$-probability}.$
@@
$\sup_{h\in BL_{1}}\left|\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}\left[h\left(\sqrt{n}(\theta-\theta_{0}\right)\right]-\mathbb{E}_{Z\sim\mathcal{N}(0,\mathcal{I}^{-1}(\theta_{0}))}[h(Z)]\right|\stackrel{{\scriptstyle P_{\theta_{0}}^{n}}}{{\rightarrow}}0,$
where $BL_{1}$ is the space of functions $h:\mathbb{R}^{p}\mapsto\mathbb{R}$ with Lipschitz norm bounded by 1. Recall that if both expectations on the left hand side were unconditional, limit of the left hand side going to 0 is the standard definition of convergence in distribution. See (van der Vaart, 2000) Sec. 23.2.1 for further discussion.
We shall henceforth continue with the $(.)\stackrel{{\scriptstyle d}}{{\rightarrow}}$ (.)  in $P_{\theta_{0}}^{n}$ probability” notation for ease of exposition.
Note that 3 allows us to apply the functional delta method (see (van der Vaart, 2000) chapter 20) to evaluate the limit of the function $\theta\mapsto\hat{p}_{k}(X_{\mathrm{test}};\theta)$ under the posterior. This yields,
$\sqrt{n}\left(\hat{p}_{k}(X_{\mathrm{test}};\theta)-\hat{p}_{k}(X_{\mathrm{test}};\theta_{0})\right)\big|\mathcal{D}_{n}\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,\sigma^{2}_{k})\qquad\textrm{in $P_{\theta_{0}}^{n}$-probability},$ (5)
where $\sigma^{2}_{k}=\left[\frac{\partial\hat{p}(X_{\mathrm{test}};\theta_{0})}{\partial\theta}^{T}\mathcal{I}^{-1}(\theta_{0})\frac{\partial\hat{p}(X_{\mathrm{test}};\theta_{0})}{\partial\theta}\right]_{k,k}$. Finally, 7 implies uniform integrability allowing us to conclude that the limit of the moment is the moment of the limiting distribution:
$\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}[\hat{p}_{k}(X_{\mathrm{test}};\theta)]\stackrel{{\scriptstyle P_{\theta_{0}}^{n}}}{{\rightarrow}}\hat{p}_{k}(X_{\mathrm{test}};\theta_{0}),$
$\mathbb{V}\textrm{ar}_{\theta\sim p(\theta|\mathcal{D}_{n})}\left[\sqrt{n}\hat{p}_{k}(X_{\mathrm{test}};\theta)\right]\stackrel{{\scriptstyle P_{\theta_{0}}^{n}}}{{\rightarrow}}\sigma_{k}^{2}.$
@@
$\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}\left[n\frac{(\hat{p}_{k}(X_{\mathrm{test}};\theta)-\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}[\hat{p}_{k}(X_{\mathrm{test}};\theta)])^{3}}{\hat{p_{k}^{\prime 2}}}\right]\stackrel{{\scriptstyle P_{\theta_{0}}^{n}}}{{\rightarrow}}0.$
@@
$n\mathrm{I}(Y_{\mathrm{test}};\theta|X_{\mathrm{test}},\mathcal{D}_{n})\stackrel{{\scriptstyle P_{\theta_{0}}^{n}}}{{\rightarrow}}\frac{1}{2}\sum_{k=1}^{K}\frac{\sigma_{k}^{2}}{\hat{p}_{k}(X_{\mathrm{test}};\theta_{0})}.$
@@ A.4 Theorem 2
Under the assumptions 1-12,
$\frac{\mathrm{I}(Y_{\mathrm{test}};\theta|X_{\mathrm{test}},\mathcal{D}_{n})}{\mathrm{I}_{b}(X_{\mathrm{test}},\mathcal{D}_{n})}\stackrel{{\scriptstyle P_{\theta_{0}}^{n}}}{{\rightarrow}}1\qquad\forall\ X_{\mathrm{test}},$
@@ A.4.1 Proof
where $\hat{p}_{k}^{\prime}$ is now some value between $\hat{p}_{k}(X_{\mathrm{test}};\hat{\theta}_{b})$ and $\mathbb{E}_{\hat{\theta}_{b}\sim p_{b}(\hat{\theta}_{b}|\mathcal{D}_{n})}[\hat{p}_{k}(X_{\mathrm{test}};\hat{\theta}_{b})]$ for k=1,,Kk=1,...,K. Next, from Theorem 10.16 of (Kosorok, 2007), we have
$\sqrt{n}\left(\hat{\theta}_{b}-\hat{\theta}\big|\mathcal{D}_{n}\right)\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,\mathcal{I}^{-1}(\theta_{0}))\qquad\textrm{in $P_{\theta_{0}}^{n}$-probability}$
@@
We can now apply the delta method for bootstrap ((van der Vaart, 2000) Theorem 23.9) to this and proceed exactly as in the Bayesian case to conclude
$n\mathrm{I}_{b}(X_{\mathrm{test}},\mathcal{D}_{n})\stackrel{{\scriptstyle P_{\theta_{0}}^{n}}}{{\rightarrow}}\frac{1}{2}\sum_{k=1}^{K}\frac{\sigma_{k}^{2}}{\hat{p}_{k}(X_{\mathrm{test}};\theta_{0})}.$
@@
$\frac{n\mathrm{I}(Y_{\mathrm{test}};\theta|X_{\mathrm{test}},\mathcal{D}_{n})}{n\mathrm{I}_{b}(X_{\mathrm{test}},\mathcal{D}_{n})}\stackrel{{\scriptstyle P_{\theta_{0}}^{n}}}{{\rightarrow}}1.$
@@ Appendix B Additional Experiments and Experimental Details
B.1 Epistemic Uncertainty Decomposition
B.1.1 CIFAR-10
@@
Experimental Details: We use the CIFAR-10 dataset (Krizhevsky, 2009) with the original training and test set folds. We use a standard ResNet-18 (He et al., 2015) architecture but modify the first convolutional layer to have a kernel size of 3, stride of 1, padding of 1 to allow the network to work with 32 $\times$32 size images. Further, we modify the last fully connected layer to only have 10 output neurons for a 10 class classification task. We use cross entropy loss and Adam (Kingma and Ba, 2015) as an optimizer with learning rate of 0.001 and betas of 0.9 and 0.999. We use a batch size of 128 and train for 15 epochs. The rest of the experiment proceeds as described in the main text.
Accuracy Metrics: Models within the deep ensemble achieve an accuracy of 0.83 on average, whereas the bootstrap models achieve an accuracy of 0.80 on average. The deep ensemble itself (i.e. averaged model predictions) achieves an accuracy of 0.89. Deep ensembles over the bootstrapped datasets (i.e. model predictions averaged over the random seeds only) achieves an accuracy of 0.88 on average across the bootstrapped datasets.
B.1.2 ImageNet
Experimental Details: We use the ImageNet dataset (Deng et al., 2009) with the original training and test set folds. For computational ease, we use a downsampled version of the images downsampled to $32\times 32$ pixels as in (Chrabaszcz et al., 2017). We use a standard ResNet-18 (He et al., 2015) architecture but modify the first convolutional layer to have a kernel size of 3, stride of 1, padding of 1 to allow the network to work with 32 $\times$32 size images. We use cross entropy loss and SGD as an optimizer with learning rate of 0.1, momentum of 0.9 and weight decay of 0.0001. We use a batch size of 128 and train for 40 epochs. We use the one cycle learning rate policy from (Smith and Topin, 2019). The rest of the experiment proceeds as described in the main text.
Accuracy Metrics: Models within the deep ensemble achieve top-1 and top-5 accuracy of 0.55 and 0.79 respectively on average, whereas the bootstrap models achieve top-1 and top-5 accuracies of 0.51 and 0.75 respectively on average. The deep ensemble itself (i.e. averaged model predictions) achieves top-1 and top-5 accuracy of 0.59 and 0.81 respectively. Deep ensembles over the bootstrapped datasets (i.e. model predictions averaged over the random seeds only) achieves top-1 and top-5 accuracy of 0.55 and 0.78 on average across the bootstrapped datasets.
B.1.3 ISIC Dataset
Experimental Details: We use the ISIC Skin Lesion challenge (2016) dataset (Gutman et al., 2016) with the original training and test set folds. We pass all images through the trained ResNet-50 from (He et al., 2015) and extract the activations from the last layer to obtain 2048-dimensional feature vectors. We then train a two layer neural network from scratch on top of these feature vectors with 512 hidden units, ReLU activation and Dropout with a rate of 0.5. We use Adam (Kingma and Ba, 2015) as an optimizer with a learning rate of 0.001 and betas of 0.9 and 0.999. We use a batch size of 32 and train for 15 epochs minimizing cross entropy loss weighted by prevalence of ground truth class prevalence in trainig data to account for class imbalance. The rest of the experiment proceeds as described in the main text.
Accuracy Metrics: Models within the deep ensemble achieve an AUC-ROC of 0.81 on average, whereas the bootstrap models achieve an accuracy of 0.79 on average. The deep ensemble itself (i.e. averaged model predictions) achieves an accuracy of 0.81. Deep ensembles over the bootstrapped datasets (i.e. model predictions averaged over the random seeds only) achieves an accuracy of 0.80 on average across the bootstrapped datasets.
B.1.4 AG News
Experimental Details: We use the AG News dataset (Zhang et al., 2015) with the original training and test set folds. We finetune the TinyBERT model from (Jiao et al., 2020). We add a linear layer on top of the final 312 dimensional output from TinyBERT to perform 4 class classification (parameters initialized randomly). We use Adam (Kingma and Ba, 2015) as an optimizer with a learning rate of 2e52e-5 and betas of 0.9 and 0.999. We also use a linear learning rate schedule with warmup as provided by the transformers (Wolf et al., 2020) library. We train for 3 epochs with a batch size of 32. We tokenize each example up to 128 tokens, padding or truncating as necessary. The rest of the experiment proceeds as described in the main text.
Accuracy Metrics: Models within the deep ensemble achieve an accuracy of 0.94 on average, whereas the bootstrap models achieve an accuracy of 0.93 on average. The deep ensemble itself (i.e. averaged model predictions) achieves an accuracy of 0.94. Deep ensembles over the bootstrapped datasets (i.e. model predictions averaged over the random seeds only) achieves an accuracy of 0.93 on average across the bootstrapped datasets.
B.2 Data Attribution